OpenStudy (anonymous):

how does sqrt(16)^(1/2) =2 ? I don't know how to get to that answer without my calculator :)

OpenStudy (anonymous):

\[(\sqrt{16})^{1/2}\] = \[\sqrt{\sqrt{16}}\] you are basically taking the sqrt two times: \[\sqrt{16} = 4\] \[4^{1/2} = \sqrt{4}\]

OpenStudy (anonymous):

\[a ^{b/c}=(\sqrt[c]{a})^{b}\]